[Brahms-l] Shift report 20050224 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Feb 24 2005 - 07:40:48 EST
Shifter leader was ebj
Shifters were 


1. RHIC delivered 5 hours of beam for this shift.  

2. I took about 1.4M trg 2 events for the FS at 2.3deg A1723.  This data does not have any T1 information.

3. I also took about 100k trg 2 events for the FS at 2.3deg A2442.  This data does contain T1 information.

4. I found that the TPC1 readout does not work in only one of the FS magents settings.  Read the elog for more details.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 24 07:42:05 2005

This archive was generated by hypermail 2.1.8 : Thu Feb 24 2005 - 07:42:19 EST